OpenStudy (anonymous):

Solve: 4(x − 1) − (x + 6) = 2

hartnn (hartnn):

4x-4-x-6=2 4x-x=2+4+6 can u solve further?

OpenStudy (anonymous):

uh no im confused

OpenStudy (anonymous):

30x-4+2+4

OpenStudy (anonymous):

30x+6-4=24

hartnn (hartnn):

from where did 30 come??

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

i added all the x's together

OpenStudy (anonymous):

4x+24x+x+x=30x

OpenStudy (anonymous):

seeee

OpenStudy (anonymous):

Hm.....

OpenStudy (anonymous):

anyone know the answer?

OpenStudy (anonymous):

im helping you get it done

hartnn (hartnn):

4x-4-x-6=2 4x-x=2+4+6 so 3x=12 so x=4

OpenStudy (anonymous):

30-2=24

OpenStudy (anonymous):



OpenStudy (anonymous):

?????_???????

OpenStudy (anonymous):

omg see im bad in math

OpenStudy (anonymous):

same xD